candy Likes pork ribs (R) and chicken (c). Her utility function is (R, C) = 10R2C. Her weekly income is k90 which she spends exclusively on R and C. The price for a slab of ribs is k10 and that for a piece of chicken wing is k15.


(a) Derive the Marshallian and Hicksian demand functions for pork ribs and chicken wings.

If the price of ribs falls to k5, algebraically and graphically show the income and substitution effects of this price change (Ribs are on the horizontal axis). Ensure to put the values on the graph showing both substitution and income effect.


1
Expert's answer
2020-09-17T08:53:01-0400
"R_1=(\frac {p_c}{2p_R})^{\frac {2}{3}}u^{\frac{1}{3}}"

"C_1=(\frac {2p_R}{p_c})^{\frac {1}{3}}u^{\frac{1}{3}}"

If "p_R=10" and "p_c=15"



"R_1=(\frac {15}{20})^{\frac {2}{3}}(20R^2C)^{\frac{1}{3}}"

"C_1=(\frac {20}{15})^{\frac {1}{3}}(20 R^2C)^{\frac{1}{3}}"



The above chart shows the initial situation in blue and green after the price change